OpenStudy (anonymous):

How do you find a limit of \[\lim{\sqrt{2+\sqrt{2 + \sqrt{2 + \sqrt{...}}}}} =\,\, ?\] ?

OpenStudy (anonymous):

is there a variable or it just limit of a const

OpenStudy (anonymous):

Yes - the variable is the number of nested operations e.g.\[ a_2= \sqrt{2+\sqrt{2}}\]

OpenStudy (anonymous):

so \[ \lim_{n\to \infty}a_n \]

OpenStudy (anonymous):

aha, so its basically testing the convergence of divergence of this sequence using limit

OpenStudy (anonymous):

use the root test to do so

OpenStudy (anonymous):

more than that : THE QUESTION IS THE LIMIT VALUE, Also I am quite scepical as to the easiness of root test application here ...

OpenStudy (anonymous):

It is surely more than 1.414 !!!!!

OpenStudy (anonymous):

It is involved, but yes.

OpenStudy (anonymous):

HEllo Messrs. @experimentX and @CliffSedge

OpenStudy (experimentx):

Uggh ... stop calling me Messrs ... I'm not a woman!!

OpenStudy (anonymous):

It looks like it converges to 2 (and rather quickly).

OpenStudy (anonymous):

LOL @experimentX He's not calling us women, he's calling us French!

OpenStudy (experimentx):

French ... Woops!! @Mikael how many language do you know??

OpenStudy (anonymous):

\[a_{n+1}=\sqrt{2+a_n}\] Recursion?

OpenStudy (anonymous):

How do you prove the convergence itself ???

OpenStudy (anonymous):

Hmm, trying to find a way to explain it . . .

OpenStudy (anonymous):

Well here is an idea 1) Show that the sequence is increasing 2) Show it is bounded by \[\sqrt{2+2}\]

OpenStudy (anonymous):

Then fundam. theorem says it converges

OpenStudy (anonymous):

2) is "un peu" trickeeee...

OpenStudy (experimentx):

Good night guys!!

OpenStudy (anonymous):

Sleep well and dream of mathless bliss !

OpenStudy (anonymous):

I guess I'm thinking of it this way: \[a_{n+1}=\sqrt{2+a_n}\] \[\rightarrow (a_{n+1})^2 = 2+a_n\] And a_n is always the same, and it began as 2. Not very rigorous, but that's how my mind sees what's happening.

OpenStudy (anonymous):

OK and the "coup de grace" is : what is the Limmit?

OpenStudy (anonymous):

You wrote it !

OpenStudy (anonymous):

IT IS CALLED FIXED POINT EQUATION\[a_n = a_{n+1}\]

OpenStudy (anonymous):

\[ LIMIT=\sqrt{2+LIMIT}\]

OpenStudy (anonymous):

Yes, it converges to 2 like I said before.

OpenStudy (anonymous):

Aha, and some courtesy to the guide .....

OpenStudy (anonymous):

My guide was MS-Excel, actually, but thanks for the fun question.

OpenStudy (anonymous):

Tahnks

OpenStudy (anonymous):

Thanks @hartnn - I am glad that we share a taste for the esthetic

hartnn (hartnn):

me too :D good question!

OpenStudy (anonymous):

I plan to create and post a generalized idea , which in mathematics is ubiquitious and very effective - the idea of \[\Large\bf\text{Fixed Point Solution}\] can you guess and suggest some directions of generalization of that ?

hartnn (hartnn):

sorry, heard that for first time.....

OpenStudy (anonymous):

Well e.g. Newton-Rapson algorithm is in fact a fixed point solution to any f(x)=0 equation using derivatives

OpenStudy (anonymous):

Here is the MAJOR IDEA: Theorem: Fixed point of Contracting Mappings in Banach space Assume that \[dist(f(x_1),f(x_2))<r*dist(x_1,x_2)\] where r=const<1

OpenStudy (anonymous):

Then the mapping f possesses a fixed point solution, and that solution is unique \[\Large f(x)= x\]

OpenStudy (anonymous):

One has to assume that the mapping f acts on a compact set K in some complete space B, and it is into that set K \[f:K\,\rightarrow K\]

OpenStudy (anonymous):

That's it

hartnn (hartnn):

i m not qualified enough to understand that....never studied such things.......sorry.

OpenStudy (anonymous):

Hey you and I are not judges - we are mutual knowledge "growers" just pointed you the idea that leads to many more places, and can be easily read upon.

hartnn (hartnn):

oh..i will surely go through them.....thanks for the references...

OpenStudy (anonymous):

That is just some very popular down-to-applications-earth application . Taken from the second reference above

OpenStudy (anonymous):

Thanx @Mikael I really enjoy when you post some thing......

OpenStudy (anonymous):

@sauravshakya it includes this topic http://www.math.ucdavis.edu/~hunter/book/ch3.pdf But MUCH BETTER TO READ FROM THE BEGINNING http://www.math.ucdavis.edu/~hunter/book/

OpenStudy (anonymous):

Thanks , friend !

OpenStudy (anonymous):

Can I ask a question?

OpenStudy (anonymous):

Ask away !

OpenStudy (anonymous):

WELL what is the solution of this: I am tired of doing this..........